### v2

@@ Abstract
Terahertz imaging holds great potential for non-destructive material inspection, but practical implementation has been limited by resolution constraints. In this study, we present a novel single-pixel THz imaging system based on a confocal microscope architecture, utilising a quantum cascade laser as both transmitter and phase-sensitive receiver. Our approach addresses these challenges by integrating laser feedback interferometry detection, achieving a two-fold improvement in lateral resolution compared to conventional reflectance confocal microscopy and a dramatic enhancement in axial resolution through precise interferometric phase measurements. This breakthrough provides lateral resolution near $\lambda/2$ and a depth of focus better than $\lambda/5$, significantly outperforming traditional confocal systems. The system can produce a 0.5 Mpixel image in under two minutes, surpassing both raster-scanning single-pixel and multipixel focal-plane array-based imagers. Coherent operation enables simultaneous amplitude and phase image acquisition, and a novel visualisation method links amplitude to image saturation and phase to hue, enhancing material characterisation. A 3D tomographic analysis of a silicon chip reveals subwavelength features, demonstrating the system’s potential for high-resolution THz imaging and material analysis. This work sets a new benchmark for THz imaging, overcoming key challenges and opening up transformative possibilities for non-destructive material inspection and characterisation.
@@ Conclusion
In conclusion, we have demonstrated the exceptional performance of a single-pixel terahertz confocal microscope employing laser feedback interferometry (LFI) for both lateral and longitudinal resolution. The system’s compact design, utilising a single quantum cascade laser for illumination and detection, simplifies alignment by integrating the roles of a detector, confocal pinhole, and illumination source within the laser itself.

### v3

@@ Abstract
Terahertz imaging holds great potential for non-destructive material inspection, but practical implementation has been limited by resolution constraints. In this study, we present a single-pixel THz imaging system based on a confocal microscope architecture, utilising a quantum cascade laser as both transmitter and phase-sensitive receiver. Our approach integrates laser feedback interferometry detection to achieve a two-fold improvement in lateral resolution and a two-order-of-magnitude enhancement in axial resolution over conventional imaging through precise interferometric phase measurements. This translates to a lateral resolution near $\lambda/2$ and a depth of focus better than $\lambda/5$, significantly outperforming traditional confocal systems. The system can produce a 0.5 Mpixel image in under two minutes, surpassing both raster-scanning single-pixel and multipixel focal-plane array-based imagers. Coherent operation enables simultaneous amplitude and phase image acquisition, and a custom visualisation method links amplitude to image saturation and phase to hue, enhancing material characterisation. A 3D tomographic analysis of a silicon chip reveals subwavelength features, demonstrating the system’s potential for high-resolution THz imaging and material analysis. This work sets a new benchmark for THz imaging, overcoming key challenges and opening up transformative possibilities for non-destructive material inspection and characterisation.
@@ Conclusion
In conclusion, we have demonstrated for the first time that LFI imaging systems provide enhanced resolution over conventional imaging, both laterally and longitudinally. The system’s compact design, utilising a single quantum cascade laser for illumination and detection, simplifies alignment by integrating the roles of a detector, confocal pinhole, and illumination source within the laser itself.